OpenStudy (anonymous):

Solve -4(3 + w) = -27 – w. Please show all of your work

OpenStudy (anonymous):

-4(3+w)=-27-w -12-4w=-27-w +27 +27 15-4w=-w +4w +4w 15=3w w=5 (when u divide 3 on both sides)